OpenStudy (anonymous):

How much would you need to invest to get $20,000 in 5 years at an annual interest rate of 8.5% compounded monthly ?

OpenStudy (campbell_st):

will using the formula \[A =P(1 +\frac{ r}{100})^n\] and the number of periods = 12 x 5 = 60 and interest rate r = 8.5/12 \[2000= P(1 + \frac{8.5 \div 12}{100})^{60}\] now solve for P

OpenStudy (anonymous):

okay i get how to solve it now..but i get 1.007 i highly dobut its right can you recheck it for me?

OpenStudy (campbell_st):

the monthly interest rate is 0.070833% or 0.00070833 so \[P = \frac{20000}{(1.00070833)^{60}}\]